OpenStudy (anonymous):

Please HELP Part 1: Simplify f(x) = 4x^2+7x-2 over 4x-1 . Part 2: Using complete sentences, explain why f(1) = 3, f(0) = 2, and f(-1) = 1, yet f(1/4) is undefined. Be sure to show your work.

OpenStudy (anonymous):

Part 1: \[x+2, (x \neq 1/4)\] Part 2: Given f(x)= (4x^2+7x-2) / 4x-1 To find f(1), you substitute 1 for x in (4x^2+7x-2) / 4x-1 Hence f(1) = (4(1)^2+7(1)-2) / (4(1)-1) =(4 + 7 - 2) / (4 - 1) = 9/3/=3 f(-1) = (4(-1)^2+7(-1)-2) / (4(-1)-1) = (4 - 7 - 2) / (-5) = -5/-5 = 1 f(1/4) = (4(1/4)^2+7(1/4)-2) / (4(1/4)-1) The denominator 4(1/4) - 1 = 1 - 1 = 0 You cannot divide by zero and therefore f(1/4) is not a real number.

OpenStudy (anonymous):

Thank You So Much! :)